what number must be added to the numerator of 3/10 to produce the fraction 4/5?
Found 2 solutions by Mathtut, stanbon:
Answer by Mathtut(3670)

first right these as like terms by finding the same denominator
:
4/5 is equal to 8/10
:
so 3/10+x=8/10
:
x=5/10
:
so the answer is 5

Answer by stanbon(48558)

what number must be added to the numerator of 3/10 to produce the fraction 4/5?
----------------------------------
Equation:
(3+x)/10 = 4/5
---
Cross-multiply to get:
5(3+x) = 10*4
15 + 5x = 40
5x = 25
x = 5
=================
Cheers,
Stan H.